# Reciprocal relation of Schwinger pair production between $\textrm{dS}_{2}$ and $\textrm{AdS}_{2}$

### v1

@@ Abstract
The Klein-Gordon and Dirac equation for a massive charged field in a uniform electric field has a symmetry of two-dimensional global de Sitter (dS) and anti-de Sitter (AdS) space. In the in-out formalism the mean numbers of spinors (spin-1/2 fermions) and scalars (spin-0 bosons) spontaneously produced by the uniform electric field are exactly found from the Bogoliubov relations both in the global and planar coordinates of (A)dS2 space. We show that the uniform electric field enhances the production of charged spinor and scalar pairs in the planar and global dS space while the AdS space reduces the pair production in which weak electric fields below the Breitenlohner-Freedman (BF) bound prohibits pair production. The leading Boltzmann factor in dS space can be written as the Gibbons-Hawking radiation or Schwinger effect with e-folding factors less than one that give the QED effect or the curvature effect. We observe that dS2 and AdS2 spaces are connected by QED, such as a reciprocal relation between the mean number of spinors and scalars provided that the spacetime curvature is analytically continued. The leading behavior of the mean numbers for spinors and scalars is explained as a sum of contour integrals of the frequency in the phase-integral formulation.
@@ I Introduction
The $D$-dimensional de Sitter (dS) and anti-de Sitter (AdS) spaces have the symmetry $O(D,1)$ and $O(D-1,2)$ respectively [1], in which the scalar or spinor fields have solutions in terms of special functions [2, 3, 4]. The symmetry has been studied in dS space [5] and AdS space [6]. The symmetry of (A)dS space gives the group theoretical representation of quantum fields [7, 8, 9]. More interestingly, the extremal limit of nonrotating charged black holes has a near-horizon geometry of AdS space [10] while the Nariai limit of nonrotating charged black holes in dS space has a near-horizon geometry of dS space [11]. Near-horizon geometries of extremal black holes are classified, and their symmetries are studied in [12].
@@
Extremal charged black holes with two coincident horizons have the zero Hawking temperature but still produce charged pairs due to an electric field on the horizon, the so-called Schwinger pair production. The Schwinger pair production of scalars has been studied in a near-horizon geometry ${\rm AdS}_{2}\times{\rm S}^{2}$ [22, 23] while catastrophic emission of charges has been found in the Nariai limit of nonrotating charged black holes in dS space, ${\rm dS}_{2}\times{\rm S}^{2}$ [24]. Rotating charged black holes have the extremal limit of warped AdS or dS space [25, 26]. Thus, the spontaneous production of charged pairs by an electric field may shed light on understanding quantum property of charged black holes [27]. The Schwinger pair production has been studied in ${\rm dS}_{2}$ [28, 29] and ${\rm AdS}_{2}$ [30].
@@
Another interesting issue of Schwinger pair production is the boundary condition imposed on quantum fields. In the vector potential for an electric field each positive frequency mode in the past scatters over a barrier and splits into another positive frequency mode and a negative frequency mode in the future, while in the Coulomb potential an incoming (incident) mode scatters by a tunneling barrier into an outgoing (reflected) mode and another outgoing (transmitted) mode [37]. It has been shown in [38, 39] that boundary conditions on quantum fields for $E_{0}/\cosh^{2}(t/T)$ and $E_{0}/\cosh^{2}(x/L)$ result in different formulas for scalars and spinors, which are inverse of each other by the correspondence: $T\leftrightarrow L$, the effective duration and distance, and $\omega\leftrightarrow k$, the energy and momentum. A similar behavior is shown for the emission formulas of charges from Nariai black holes in the timelike and spacelike regions [24].
@@
In this paper we will study the Schwinger pair production of spinors (spin-1/2 fermions) and scalars (spin-0 bosons) by a uniform electric field in the global and planar coordinates of ${\rm(A)dS}_{2}$ and investigate the relation between the mean numbers in both coordinates. In the planar coordinates of ${\rm(A)dS}_{2}$ the Schwinger pair production has been studied for scalars [28, 30, 29, 40] and spinors [41, 42, 43, 44, 45], and in the global coordinates the scalar production has been studied in [46]. However, to the best of our knowledge, the Schwinger pair production of spinor has not been studied in the global coordinates. As the uniform electric field does not break the symmetry of ${\rm(A)dS}_{2}$ in two dimensions, the Klein-Gordon equation for a massive charged scalar and the Dirac equation for a massive charged spinor can be solved in terms of hypergeometric functions or Whittaker functions, which lead to the exact formulas for Schwinger pairs produced by the electric field which is assisted or suppressed by the scalar curvature. The spinor or scalar production increases in dS space while it decreases in AdS space. There is no bosonic amplification for scalars and the Pauli blocking suppresses the mean number less than unity, though more spinors are produced than scalars in AdS space.
@@
Because of the symmetry the mean numbers can be expressed in terms of the Maxwell (invariant) scalar for the electric field and the spacetime curvature, which have the correct limits of the Minkowski space and pure dS space without an electric field, and whose leading Boltzmann factor is given by the ratio of the Maxwell scalar to the scalar curvature square. The exact Schwinger formulas may shed light on understanding how the Maxwell field strongly couples to the spacetime curvature and vice versa. Remarkably, the mean numbers for spinors shows the reciprocal relation: the mean number in the dS space is the inverse of that in AdS space provided that the curvature is analytically continued. We apply the phase-integral formulation to explain a physical origin of the leading Boltzmann factor as contributions from simple poles in a complex plane of time (space).
@@
The organization of this paper is as follows. In section II we study the Schwinger pair production for scalars and spinors produced by a uniform electric field in the global coordinates of ${\rm dS}_{2}$. In section III the mean numbers for scalars and spinors produced by a uniform electric field are found in the global coordinates of ${\rm AdS}_{2}$. In section IV we find the mean number of spinor production in the planar coordinates of ${\rm(A)dS}_{2}$. In section V we discuss the reciprocal relation between the mean numbers both in the global ${\rm dS}_{2}$ and ${\rm AdS}_{2}$ and in the planar ${\rm dS}_{2}$ and ${\rm AdS}_{2}$. In section VI we give a physical interpretation of the Schwinger formulas in the phase-integral formulation. In section VII we discuss the physics of the Schwinger production of spinors and scalars.
@@ II Pair Production in Global de Sitter Space
In finding explicit formulas for the Schwinger pair production in any dimensional global dS space, an external uniform electric field prohibits both Klein-Gordon and Dirac equations from separating into variables, except for dS2 with the scalar curvature $R=2H^{2}$:
@@
The electric field points the positive (counterclockwise) direction of $\theta$. The Maxwell tensor has components $F_{01}=-F_{10}=-E\cosh(Ht)/H$ and gives the Maxwell scalar ${\cal F}=-F_{\mu\nu}F^{\mu\nu}/4=E^{2}/2$.
@@ II.1 Klein-Gordon and Dirac equations
after the space integration, in terms of positive (+)(+)( + ) and negative ()(-)( - ) frequency modes becomes
@@ II.2 Charged Scalar Production in dS2
The equation is invariant under time reversal $\tau\to-\tau$ together either $\ell\to-\ell$ or $q\to-q$. The general solution is given by hypergeometric functions:
@@
For the in-state (adiabatic vacuum) in the past infinity ($\tau=Ht\to-\infty$) in a Riemann sheet where
@@
Therefore, the corresponding positive (+)(+)( + ) and negative ()(-)( - ) frequency solutions are
@@
We impose the boundary condition $\phi_{\mathrm{in}}^{(-)}=0$ that describes the scattering of a positive frequency mode of the in-state into a positive and a negative frequency mode of the out-state:
@@
and the in- and out-states reduce to
$\phi_{\mathrm{in}}^{(+)}$ $=$ $2^{1+\ell-i2\mu}\,C_{1}B\,\mathrm{e}^{i\pi(1+2\ell)/4}\,\mathrm{e}^{\pi(\kappa-\mu)/2}\Bigl{(}1-\mathrm{e}^{2\pi\mu}|A|^{2}/|B|^{2}\Bigr{)}\mathrm{e}^{-\tau/2-i\mu\tau},$
@@
Note that the broken time-reversal symmetry $\tau\to-\tau$ due to the electric field in (9) is restored by the charge conjugation $q\to-q$, which implies that the positive frequency of particle in the future corresponds to that of antiparticle in the past, and vice versa, and whose property is not shared by the planar coordinates.
The Bogoliubov coefficients $\alpha$ and $\beta$ defined as [24]
@@
It should be noted that the mean number (24) is invariant under charge conjugation and independent of the mode $\ell$ or its direction. It is a consequence of the mean number ${\cal N}=\langle\textrm{in}|N_{\textrm{out}}|\textrm{in}\rangle=\langle\textrm{out}|N_{\textrm{in}}|\textrm{out}\rangle$ and the equality of produced particles and antiparticles, which are invariant under both the charge conjugation and time reversal. The vacuum persistence amplitude square
@@
Note that the leading Boltzmann factor is ${\cal N}^{\rm(sc)}_{\rm dS}\approx\mathrm{e}^{-2\pi(\mu-\kappa)}$ and reduces to the Schwinger formula ${\cal N}=\mathrm{e}^{-\pi m^{2}/qE}$ in the flat spacetime limit $(H=0)$.
@@ II.3 Charged Spinor Production in dS2
which can be decoupled to two components in the second-order formulation
@@
A physically interesting interpretation of the mean number is the effective temperature. Two temperatures are associated with the electric field and dS space:
@@
The leading Boltzmann factor is ${\cal N}^{\rm(sp)}_{\rm dS}\simeq e^{-(m/T_{\rm GH})\times((\sqrt{T_{\rm U}^{2}+T_{\rm GH}^{2}}-T_{\rm U})/T_{\rm GH})}$. In the zero electric field, namely $T_{\rm U}=0$ and $T_{\rm eff}=T_{\rm GH}$, the mean number becomes Gibbons-Hawking radiation
@@
Moreover, when $T_{\rm H}\gg T_{\rm GH}$ in which the dS radius is larger than the radius of circular motion in Euclidean time ($l_{\rm dS}=1/H\gg m/qE$), $T_{\rm eff}\simeq 2T_{\rm U}+T_{\rm GH}^{2}/2T_{\rm U}$ and the mean number becomes approximately ${\cal N}\simeq\mathrm{e}^{-m/2T_{\rm U}}$, the Schwinger formula in the Minkowski space. In the other case of $T_{\rm GH}\gg T_{\rm U}$, the Boltzmann factor becomes ${\cal N}\simeq\mathrm{e}^{-m/T_{\rm GH}\times(1-T_{\rm U}/T_{\rm GH})}$. The emission of charges enhanced by the electric field may be interpreted as Gibbons-Hawking radiation of massive charge with a chemical potential $qE/H=qEl_{\rm dS}$.
@@
${\cal N}^{\rm(sp)}_{\rm dS}=\begin{cases}\mathrm{e}^{-\frac{m}{T_{\rm GH}}\times\frac{1}{\sqrt{1+(T_{\rm U}/T_{\rm GH})^{2}}+T_{\rm U}/T_{\rm GH}}}\quad(T_{\rm GH}>T_{\rm U}),\\ \mathrm{e}^{-\frac{m}{2T_{\rm U}}\times\frac{2}{\sqrt{1+(T_{\rm GH}/T_{\rm U})^{2}}+1}}\quad(T_{\rm U}>T_{\rm GH}).\end{cases}$ (47)

### v2

@@ Abstract
The Klein-Gordon and Dirac equation for a massive charged field in a uniform electric field has a symmetry of two-dimensional global de Sitter (dS) and anti-de Sitter (AdS) space. In the in-out formalism the mean numbers of spinors (spin-1/2 fermions) and scalars (spin-0 bosons) spontaneously produced by the uniform electric field are exactly found from the Bogoliubov relations both in the global and planar coordinates of (A)dS2 space. We show that the uniform electric field enhances the production of charged spinor and scalar pairs in the planar and global dS space while the AdS space reduces the pair production in which weak electric fields below the Breitenlohner-Freedman (BF) bound prohibits pair production. The leading Boltzmann factor in dS space can be written as the Gibbons-Hawking radiation or Schwinger effect enhanced by e-folding factors less than one that give the QED effect or the curvature effect. We observe that dS2 and AdS2 spaces are connected by QED, such as a reciprocal relation between the mean number of spinors and scalars provided that the spacetime curvature is analytically continued. The leading behavior of the mean numbers for spinors and scalars is explained as a residue sum of contour integrals of the frequency or momentum in the phase-integral formulation.
@@ I Introduction
The $D$-dimensional de Sitter (dS) and anti-de Sitter (AdS) spaces have the symmetry $O(D,1)$ and O(D1,2)O(D-1,2) respectively [1], in which the scalar or spinor fields have solutions in terms of special functions [2, 3, 4]. The symmetry has been studied in dS space [5] and AdS space [6]. The symmetry of (A)dS space gives the group theoretical representation of quantum fields [7, 8, 9]. More interestingly, the extremal limit of nonrotating charged black holes has a near-horizon geometry of AdS space [10] while the Nariai limit of nonrotating charged black holes in dS space has a near-horizon geometry of dS space [11]. Near-horizon geometries of extremal black holes are classified, and their symmetries are studied in [12].
@@
Extremal charged black holes with two coincident horizons have the zero Hawking temperature but still produce charged pairs due to an electric field on the horizon, the so-called Schwinger pair production. The Schwinger pair production of scalars has been studied in a near-horizon geometry ${\rm AdS}_{2}\times{\rm S}^{2}$ [22, 23] while catastrophic emission of charges has been found in the Nariai limit of nonrotating charged black holes in dS space, ${\rm dS}_{2}\times{\rm S}^{2}$ [24]. Rotating charged black holes have the extremal limit of warped AdS or dS space [25, 26]. Thus, the spontaneous production of charged pairs by an electric field may shed light on understanding quantum property of charged black holes [27]. The Schwinger pair production of scalars has been studied in ${\rm dS}_{2}$ [28, 29] and ${\rm AdS}_{2}$ [30].
@@
Another interesting issue of Schwinger pair production is the boundary condition imposed on quantum fields. In the vector potential for an electric field each positive frequency mode in the past scatters over a barrier and splits into another positive frequency mode and a negative frequency mode in the future, while in the Coulomb potential an incoming (incident) mode scatters by a tunneling barrier into an outgoing (reflected) mode and another outgoing (transmitted) mode [37]. It has been shown in [38, 39] that boundary conditions on quantum fields for $E_{0}/\cosh^{2}(t/T)$ and $E_{0}/\cosh^{2}(x/L)$ result in different formulas for scalars and spinors, which are inverse of each other by the correspondence: $T\leftrightarrow L$, the effective duration and distance, and $\omega\leftrightarrow k$, the energy and momentum. A similar behavior is shown for the emission formulas of scalars from extremal RN and Nariai black holes [23, 24].
@@
In this paper, we present a comprehensive study of Schwinger pair production for spinors (spin-1/2 fermions) and scalars (spin-0 bosons) induced by a uniform electric field in both the global and planar coordinates of (A)dS2. We also investigate the relationship between the mean numbers of pair production obtained in both coordinates of these two geometries. In the planar coordinates of (A)dS2, Schwinger pair production has been previously studied for scalars [28, 30, 29, 40] and spinors [41, 42, 43, 44, 45]. Scalar pair production in the global coordinates was examined in [46]. However, to the best of our knowledge, the Schwinger pair production of spinors in the global coordinates of both dS2 and AdS2 spaces has not yet been investigated. As the uniform electric field does not break the symmetry of ${\rm(A)dS}_{2}$ in two dimensions, the Klein-Gordon equation for a massive charged scalar and the Dirac equation for a massive charged spinor can be solved in terms of hypergeometric functions or Whittaker functions, which lead to the exact formulas for Schwinger pairs produced by the electric field which is assisted or suppressed by the scalar curvature. The spinor or scalar production increases in dS space while it decreases in AdS space. The bosonic amplification for scalars is not found, and the Pauli blocking suppresses the mean number less than unity, though more spinors are produced than scalars in AdS space.
@@
Because of the symmetry the mean numbers can be expressed in terms of the Maxwell (invariant) scalar for the electric field and the spacetime curvature, which have the correct limits of the Minkowski space and pure dS space without an electric field, and whose leading Boltzmann factor is given by the ratio of the Maxwell scalar to the scalar curvature square. The exact Schwinger formulas may shed light on understanding how the Maxwell field strongly couples to the spacetime curvature and vice versa. Remarkably, our results reveal that the mean number of produced scalars or spinors satisfies a reciprocal relation: upon analytic continuation of the curvature, the mean number in dS space becomes the inverse of that in AdS space and vice versa. [^1] The reciprocality of Schwinger effect between the dS and AdS space in this paper and also the spatial and temporal Sauter-type electric fields should not be confused with the reciprocity relations in thermodynamics, such as the Maxwell and the Onsager-Casimir reciprocity relations [47]. We apply the phase-integral formulation to explain a physical origin of the leading Boltzmann factor as contributions from simple poles of frequency or momentum in a complex plane of time (space).
@@
The organization of this paper is as follows. In section II we study the Schwinger pair production for scalars and spinors produced by a uniform electric field in the global coordinates of ${\rm dS}_{2}$. In section III the mean numbers for scalars and spinors produced by a uniform electric field are found in the global coordinates of ${\rm AdS}_{2}$. In section IV we find the mean number of spinor production in the planar coordinates of ${\rm(A)dS}_{2}$. In section V we discuss the reciprocal relation between the mean numbers of scalars and spinors both in the global ${\rm dS}_{2}$ and ${\rm AdS}_{2}$ and in the planar ${\rm dS}_{2}$ and ${\rm AdS}_{2}$. In section VI we give a physical interpretation of the Schwinger formulas in the phase-integral formulation. In section VII we discuss the physics of the Schwinger production of spinors and scalars.
@@ II Pair Production in Global de Sitter Space
In finding explicit formulas for the Schwinger pair production in any dimensional global dS space, an external uniform electric field prohibits both Klein-Gordon and Dirac equations from separating by spherical harmonics, except for dS2 with the scalar curvature $R=2H^{2}$:
@@
The electric field points the positive (counterclockwise) direction of $\theta$. The Maxwell tensor has components $F_{01}=-F_{10}=-E\cosh(Ht)/H$ and gives the Maxwell scalar ${\cal F}=-F_{\mu\nu}F^{\mu\nu}/4=E^{2}/2$. The time reversal symmetry of the global metric (1) and the anti-symmetry of the gauge field (2) lead to the time reversal symmetry of the Klein-Gordon or Dirac equation with minimal coupling under the charge conjugation, $q\rightarrow-q$, which is not shared by the planar coordinates, as will be shown explicitly below. It is the two-dimensional $CPT$ invariance of the field equation.
@@ II.1 Klein-Gordon and Dirac equations
after the space integration, in terms of positive $(+)$ and negative ()(-) frequency modes becomes
@@ II.2 Charged Scalar Production in dS2
As the global dS metric is time reversal symmetric, the equation is invariant under $\tau\to-\tau$, together either $\ell\to$-$\ell$ or $q\to-q$. Once we find a solution, this symmetry can help one get the second solution, for example by $\tau\to-\tau,q\to-q$. The general solution is given by hypergeometric functions:
@@
Following Refs. [48] for scalar and [49] for spinor, the in-state (asymptotically WKB (adiabatic) vacuum) in the past infinity ($\tau=Ht\to$-$\infty$) in a Riemann sheet where
@@
Therefore, the corresponding positive $(+)$ and negative ()(-) frequency solutions are
@@
We impose the boundary condition $\phi_{\mathrm{in}}^{(-)}=0$ that describes the scattering of a positive frequency mode of the in-state into a positive and a negative frequency mode of the out-state [^2] As will be shown in Sec. VI, the canonical time-dependent form (97) for a spinless scalar describes a scattering process with energy $\mu^{2}$ over a negative potential barrier.
@@
then the in- and out-states reduce to
$\phi_{\mathrm{in}}^{(+)}$ $=$ $2^{1+\ell-i2\mu}\,C_{1}B\,\mathrm{e}^{i\pi(1+2\ell)/4}\,\mathrm{e}^{\pi(\kappa-\mu)/2}\Bigl(1-\mathrm{e}^{2\pi\mu}|A|^{2}/|B|^{2}\Bigr)\mathrm{e}^{\tau/2-i\mu\tau},$
@@
Note that the broken time reversal symmetry $\tau\to-\tau$ due to the electric field in (9) is restored by the charge conjugation $q\to-q$, which implies that the positive frequency of particle in the future corresponds to that of antiparticle in the past, and vice versa, and whose property is not shared by the planar coordinates since the metric covers different patch under flipping of time coordinate.
For a scattering process, the Bogoliubov coefficients $\alpha$ and $\beta$ defined as [24], with normalized $\phi_{\mathrm{in}}^{(+)},\phi_{\mathrm{out}}^{(+)}$ and $\phi_{\mathrm{out}}^{(-)}$ without causing confusion of notation
@@
It should be noted that the mean number (24) is invariant under charge conjugation and independent of the mode $\ell$ or its direction. The invariance under charge conjugation is a consequence of the mean number ${\cal N}=\langle\textrm{in}|N_{\textrm{out}}|\textrm{in}\rangle=\langle\textrm{out}|N_{\textrm{in}}|\textrm{out}\rangle$ and the equality of produced particles and antiparticles, which are invariant under both the charge conjugation and time reversal of the Klein-Gordon or Dirac equation in the global coordinates. The independence of $\ell$ may be understood from the infinite redshift of physical momentum $\ell_{\rm phys}=\ell/(\cosh(Ht)/H)$ at $t=\pm\infty$, which can be seen by writing the $\tau$-dependent frequency square in (9) as $(qE/H^{2}\tanh\tau+\ell/\cosh\tau)^{2}+(m/H)^{2}$. Similarly, the particle production from the global coordinates is independent of the harmonic numbers [31], which is a zero-field limit $E=0$ of (24) in two dimensions. The vacuum persistence amplitude square
@@
Note that the leading Boltzmann factor is ${\cal N}^{\rm(sc)}_{\rm dS}\approx\mathrm{e}^{-2\pi(\mu-\kappa)}$ and reduces to the Schwinger formula ${\cal N}=\mathrm{e}^{-\pi m^{2}/qE}$ in the flat spacetime limit $(H=0)$. Interestingly, there is no bosonic amplification from the dS space.
@@ II.3 Charged Spinor Production in dS2
whose components decouple from each other in the second-order formulation
@@
A physically interesting interpretation of the mean number is the effective temperature [40, 50, 51]. Two temperatures are associated with the electric field and dS space:
@@
It should be noted that the Unruh temperature in dS space is $\sqrt{T_{\rm U}^{2}+T_{\rm GH}^{2}}$ [52] and that the effective temperature (45) is a characteristic of the Schwinger effect, which becomes $T_{\rm eff}=2T_{\rm U}$ in the Minkowski space. When $mT_{\rm U}\geq 2\pi T_{\rm GH}^{2}$, i.e, the dS radius is larger than the radius of the local circular orbit of charge in Euclidean time ($l_{\rm dS}=1/H\geq l_{\rm E}=1/\sqrt{qE}$), the mean number is approximated by the leading Boltzmann factor ${\cal N}^{\rm(sp)}_{\rm dS}\simeq\mathrm{e}^{-(m/T_{\rm GH})\times(\sqrt{T_{\rm U}^{2}+T_{\rm GH}^{2}}-T_{\rm U})/T_{\rm GH}}$. In the zero electric field, namely $T_{\rm U}=0$ and $T_{\rm eff}=T_{\rm GH}$, the mean number becomes Gibbons-Hawking radiation
@@
Moreover, when $T_{\rm U}\gg T_{\rm GH}$ ($l_{\rm dS}=1/H\gg m/qE$), $T_{\rm eff}\simeq 2T_{\rm U}+T_{\rm GH}^{2}/2T_{\rm U}$ and the mean number becomes approximately ${\cal N}\simeq\mathrm{e}^{-m/2T_{\rm U}}$, the Schwinger formula in the Minkowski space. In the other case of $T_{\rm GH}\gg T_{\rm U}$, the Boltzmann factor becomes ${\cal N}\simeq\mathrm{e}^{-m/T_{\rm GH}\times(1-T_{\rm U}/T_{\rm GH})}$. The emission of charges enhanced by the electric field may be interpreted as Gibbons-Hawking radiation of massive charge with a chemical potential $qE/H=qEl_{\rm dS}$.
@@
${\cal N}^{\rm(sp)}_{\rm dS}=\begin{cases}\mathrm{e}^{-\frac{m}{T_{\rm GH}}\times\frac{1}{\sqrt{1+(T_{\rm U}/T_{\rm GH})^{2}}+T_{\rm U}/T_{\rm GH}}},\quad&(T_{\rm GH}>T_{\rm U}),\\ \mathrm{e}^{-\frac{m}{2T_{\rm U}}\times\frac{2}{\sqrt{1+(T_{\rm GH}/T_{\rm U})^{2}}+1}},\quad&(T_{\rm U}>T_{\rm GH}).\end{cases}$ (47)